Max went to the fair with 5 friends. Some paid $2 each to ride the bumper cars. The rest paid $3 each to ride the roller coaster. They spent $16 in all. How many rode each ride?
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16949)

let the bumper car ride cost x
Let roller coaster ride cost y
Total money spent = 16
2x+3y=6
Everybody has sat on a ride
x+y =6
y=6-x
2x+3y=16
plug the value of y in the equation
2x+3(6-x)=16
2x+18-3x=16
-x=16-18
-x = -2
x=2 bumper car rides
spo y =4 roller coaster rides

Answer by richwmiller(17219)

x+y=5
x=5-y
2(5-y)+3y=16
Double check your numbers.
Let's say all of them got on the bumper cars 2*5=10
3*5=15
Even if all 5 got on the $3 ride they would not spend $16.
It would work if some got on both. But you don't say that.